Sunday, September 5, 2010

Nightlog 2010-09-04

SA: Paul
SO: Patrick
Others: Zolisa, Siphelo, Hitesh


- We did some more out of focus tests to see if the get the same variation =
flux distribution over the VI stage.

- We found that that auto collimator has been very stable and we detected no
drift which could have been one of the explanations for the observed images=

- Hitesh came out to have a look at the images we were getting and agrees t=
all is not right. We are speculating that its a mis-aligned VI stage. This
might have occurred when it was put back after the rho ring measurements. We
will only know for sure when the guys come up tomorrow night to check the VI
stage alignment.

- We closed early, we can't do any IQ with the current aberrations. =C2=A0=


No comments:

Post a Comment